# ESGSenticNet: A Neurosymbolic Knowledge Base for Corporate Sustainability Analysis

### v1

@@ 1 Introduction
With growing concerns over social inequality and global warming, it is imperative to ensure the sustainability of public companies [13]. To this end, accurate insights into the sustainability efforts of corporations are crucial to drive responsible business practices [40]. However, the complexity and volume of corporate sustainability data [38], often disclosed through sustainability reports (i.e. sustainability disclosures), can make these insights inaccessible without specialised tools.
While NLP tools can unlock insights from this vast array of corporate sustainability data, their deployment remains nascent within this field, often lacking the specialised design needed for corporate sustainability analysis. Specifically, they do not address the challenges of immateriality, complexity and subjectivity, which are frequently associated with sustainability data [41].
  • Immateriality: Sustainability disclosures are lengthy, containing significant amounts of irrelevant content for sustainability analysis [51, 28]. Key or material information for stakeholders can often be obfuscated [49], making their extraction via NLP tools challenging.
  • Complexity: Sustainability ideas are sophisticated, and are typically expressed by complex syntactic constructions [50]. The distinctiveness in sustainability language necessitates that NLP tools must encompass more than just general language understanding capabilities.
  • Subjectivity: Stakeholders are interested in diverse sustainability areas, and utilise different frameworks and methods for sustainability analysis [39, 46]. Therefore, designing an NLP tool that comprehensively meets the needs of different stakeholders is challenging.
As a result, NLP algorithms, including the state of the art, are largely ineffective at extracting useful corporate sustainability insights. In our experiments later in this paper (section 6.3), we validate this ineffectiveness by benchmarking these algorithms in the sustainability topic analysis task.
In response to these limitations, we develop ESGSenticNet, a publicly available knowledge base that comprises 44k sustainability knowledge triplets from 23k unique concepts, specifically designed to mine insights from sustainability disclosures. To address the aforementioned issues of complexity, immateriality and subjectivity, ESGSenticNet is constructed from a neurosymbolic framework that blends symbolic linguistics rules with sub-symbolic reasoning through deep learning. The symbolic component of this framework involves specialised concept parsing that leverages predefined linguistic patterns and rules within the sustainability domain to identify meaningful concepts within sustainability text, addressing the complexity of sustainability language. Additionally, these rules focus on the most pertinent linguistic structures in sustainability, enabling the identification of key sustainability information such as sustainability efforts and actions to reduce immateriality [7]. Building on the structured outputs of the symbolic component, the subsymbolic component integrates GPT-4o inference with semi-supervised label propagation to classify the relation between the parsed concepts and their respective sustainability categories. By interpreting nuanced language variations and contextual subtleties that symbolic rules cannot fully capture, this component ensures that concepts are organised according to key sustainability categories, further enhancing their materiality. Additionally, these categories within ESGSenticNet are arranged according to a hierarchical and comprehensive taxonomy, addressing subjectivity by allowing concepts to be leveraged at different levels of granularity, tailored to the specific interest of stakeholders. The result is an extensive knowledge base, ESGSenticNet, that comprises actionable information in the form of knowledge triplets. These triplets highlight concepts and their relation toward sustainability themes, such as (‘minimise paper waste’ (concept), supports (relation), ‘waste management’ (category)).
We evaluate ESGSenticNet’s ability to extract meaningful insights from sustainability disclosures, by deploying the knowledge base as a structured lexicon method for sustainability topic analysis. In this task, the concepts within ESGSenticNet’s knowledge triplets are matched from a large sustainability corpora. ESGSenticNet improves on established baselines significantly, more extensively capturing sustainability related information, including material content that highlight sustainability efforts. ESGSenticNet yields a significant number of ESG topic terms, outperforming existing state of the art baselines by at least 26% and 31% respectively on the ESG relatedness and ESG action orientation of these terms. This highlights demonstrates ESGSenticNet’s potential to contribute meaningfully to sustainability analysis. Moreover, as a lexical method, ESGSenticNet is able to extract ESG topic terms without any training or significant technical expertise. This significantly allows it to be leveraged by stakeholders of different backgrounds.
The rest of our paper is arranged as follows. We first develop a comprehensive and hierarchical sustainability taxonomy for organising our concepts (section 4). Secondly, we delve into the extraction of concepts via ESG Concept Parser (section 5.2), processing the concepts to ensure coherence (section 5.3), as well as our methodology for classifying these concepts via GPT-4o and semi-supervised label propagation (sections 5.45.55.65.7). Finally, we evaluate the accuracy of ESGSenticNet labels (section 5.9), before running extensive experiments to assess the effectiveness of ESGSenticNet for sustainability analysis against established baselines (section 6). The development process for ESGSenticNet is summarised by figure 1, showcasing the key steps for concept parsing and taxonomy derivation, and how they are intertwined with the labelling process (1-4).
2 Literature Review
NLP-driven Sustainability Analysis. To enhance sustainability analysis, NLP can be deployed for different tasks, such as analysing sustainability topics [65], sustainability sentiment classification [43], extracting textual patterns [8], among other tasks. The insights extracted from NLP can powerfully reduce time costs and increase the consistency of evaluation [2, 15], making NLP a useful tool for sustainability analysis. Yet, despite NLP’s potential applications, it has only been explored within sustainability analysis to a limited degree, often lacking specialised design [41]. This motivates the development of specially designed NLP tools for corporate sustainability analysis.
Lexicons for Sustainability Analysis. For corporate sustainability analysis, lexical approaches have been widely leveraged for textual analysis [22], due to their simplicity and relevance to financial text analysis [53, 55]. To facilitate lexical analysis, sustainablity-specific lexicons have been developed. Approaches such as Baier’s dictionary [6], have focused on developing single-word lexicons according to ESG pillars. However, while Baier’s dictionary has been a popular choice for sustainability analysis works [25], single word lexicons cannot adequately capture the semantics of the sustainability field. In contrast, dictionaries that have collected sustainability related text spans (of multiple-grams) [17, 29], are limited, not publicly available, and not comprehensively labelled beyond ESG pillars. Additionally, due to their reliance on manual human-annotation, these dictionaries have an exceedingly small number of lexicons (less than 800), raising questions on their comprehensiveness, and consequently their effectiveness for capturing textual patterns. This motivates the development of an extensive knowledge base that can function as a structured dictionary of concepts, and the exploration of automated methods such as NLP for constructing and organising sustainability lexicons.
Knowledge Base Construction. Methods for constructing knowledge bases can vary according to the domain and application [26]. Notable knowledge bases, such as ConceptNet [31] involved concept parsing and human annotation, SenticNet 5 [11] leveraged deep learning and lexical substitution, question answering via advanced graphical neural networks (GNNs) [63]. Newer avenues for knowledge construction involve generative-LLMs, which have shown strong performance in related tasks such as reasoning, construction, and generalisation [67]. Unlike all the previous methods which heavily rely on semantic similarities for labelling, generative LLMs possess powerful emergent capabilities that extend beyond textual similarities [58]. These include reasoning, knowledge relationship construction, and generalisation to different domains [67]. Moreover, unlike GNNs that rely on heavily structed text data [32], generative-LLMs can flexibly adapt to unstructured text. This, coupled with added domain flexibility and inferencing capabilities on limited textual cues, motivates our use of generative-LLMs for knowledge derivation within sustainability.
3 Data
1680 sustainability disclosures from a wide range of Singapore Exchange (SGX) based companies, in the period of 2017 to 2022 inclusive, are utilised for taxonomy and knowledge base construction. These reports can be found online via respective company websites, and on request, we can provide a list of the reports utilised (i.e. the companies and year of reporting).
Figure 1: Overview of ESGSenticNet Development
4 Sustainability Taxonomy
[…]

### v2

@@ 1 Introduction
Amid growing demands for corporate sustainability—companies’ commitment to reducing environmental impact, promoting social responsibility, and upholding ethical governance [54]—regulators and investors are increasingly focused on corporate sustainability analysis [16]. This involves evaluating a company’s progress toward Environmental, Social, Governance (ESG) goals, otherwise known as ESG assessment, with company-disclosed sustainability reports often used for analysis [52]. However, the complexity and sheer volume [48] of these reports typically render manual analysis impractical [53]. To overcome this challenge, NLP-driven ESG topic analysis — the automatic identification of ESG topics within sustainability reports — has been widely implemented. By enabling the automated and structured analysis of sustainability reports, ESG topic analysis provides regulators and investors key insights on a corporation’s sustainability performance, such as extent of alignment with environmental objectives [70], future orientation of sustainability activities [34], among others. These insights are useful for a variety of applications, such as sustainable portfolio construction [35] and regulatory auditing [27], which significantly underpin the field of sustainable finance.
However, despite advancements in NLP, existing methods struggle to meet the specialised demands of ESG topic analysis, limiting their effectiveness for this task. ESG topic analysis often aims to capture ESG-related content that is both relevant for ESG assessments and aligned with established ESG frameworks [54], such as SASB [25] and GRI [43]. To accomplish this, commonly used NLP techniques, including rule-based dictionary methods [5] and unsupervised topic modeling [10], may fall short as they lack task-specific design. We identify three key shortcomings below:
  • Irrelevance: Existing topic modeling methods, while excellent at producing topic terms without supervision [1], cannot be tuned to capture ESG content. Therefore, captured topic terms may not be relevant to ESG, leading to topics irrelevant for ESG analysis. Additionally, sustainability dictionaries that are popularly adapted for ESG topic analysis - i.e Baier [37], comprise a considerable portion of non-ESG terms.
  • Immateriality: Stakeholders often prioritise content that explicitly conveys the ESG efforts or actions undertaken by corporations, as this content has greater materiality or significance for ESG performance [8]. However, both rule-based dictionaries and topic models do not focus on extracting topic terms that capture ESG-oriented actions. To better align with stakeholder interests, topic terms should be refined to reflect these actions.
  • Limited Organisation: Due to its unsupervised nature, topic modeling may generate topics that are not organised according to ESG frameworks. Consequently, even if ESG-related topics emerge, their lack of organisation means that they fail to comprehensively capture all dimensions of ESG. This incomplete coverage can be problematic for sustainability stakeholders, who rely on structured ESG frameworks to ensure that all relevant ESG aspects are analysed [44, 26]. While sustainability dictionaries offer a more organised approach [34], their limited size (fewer than 800 terms) constrains their coverage, potentially restricting the range of ESG topics identified in practice.
Addressing the above limitations requires an ESG topic analysis approach that explicitly captures relevant, material, and well-structured ESG content. To this end, we explore the following research questions:
  • RQ1: How can a topic analysis method be designed to mitigate the limitations of existing NLP techniques in capturing relevant and material ESG content according to structured ESG frameworks?
  • RQ2: What criteria and evaluation methods are appropriate for assessing the effectiveness of ESG topic analysis?
  • RQ3: To what extent does a method specifically designed for ESG topic analysis outperform existing NLP approaches for this task?
To contribute to these research questions, we develop ESGSenticNet, an NLP tool specifically designed to facilitate ESG topic analysis. We then evaluate it extensively against state-of-the-art NLP methods, using specialised evaluation metrics tailored for ESG topic analysis.
ESGSenticNet is a publicly available knowledge base comprising 44k sustainability knowledge triples from 23k unique concepts. While ESGSenticNet itself is a knowledge base, it functions as a structured lexicon for rule-based topic analysis. To elaborate, each triple within ESGSenticNet follows the format (concept, relation, ESG topic)— i.e. (‘minimise paper waste’, supports, ‘waste management’). During topic analysis, the concepts serve as topic terms, which can be matched within sustainability reports to indicate the presence of specific ESG topics. Unlike existing topic analysis methods, ESGSenticNet explicitly addresses irrelevance, immateriality and limited organisation, through its construction from a neurosymbolic framework. This framework integrates symbolic linguistic rules and sub-symbolic deep-learning to ensure ESGSenticNet’s triples comprise relevant and material ESG concepts that are effectively organised according to structured ESG topics.
To reduce irrelevance, the symbolic component applies sustainability-specific linguistic patterns to extract ESG-related concepts, thereby minimising the occurrence of irrelevant topic terms in ESG topic analysis. To address immateriality, the applied linguistic patterns also prioritise extracting concepts that explicitly describe ESG actions, making topic terms more meaningful for stakeholders seeking actionable ESG insights. To ensure organisation, the sub-symbolic component leverages GPT-4o inference, supported with a graph-based semi-supervised method, to categorise concepts according to structured ESG topics under our constructed taxonomy. Put together, the neurosymbolic framework yields triples in the form of (concept, relation, ESG topic), and ensures that ESGSenticNet addresses the shortcomings of existing NLP for ESG topic analysis.
We evaluate ESGSenticNet’s effectiveness in ESG topic analysis by benchmarking it against state-of-the-art NLP algorithms commonly used for this task. Given the unique requirements of ESG topic analysis—capturing ESG content material for assessments, and aligned with established ESG frameworks—we introduce specialised evaluation metrics, namely (1) ESG-relatedness: The proportion of all topic terms relevant to ESG; (2) ESG-action orientation: The proportion of all topic terms focused on ESG actions, representing their material significance in sustainability assessments; (3) Unique ESG topic terms: The number of unique ESG terms, which highlights the comprehensiveness and diversity of ESG content captured. (1) and (2) directly addresses the core challenges of irrelevance and immateriality respectively, and (3) is related to the method’s practical utility in real-world deployments.
We show that ESGSenticNet significantly outperforms existing baselines on these three metrics, achieving 76% ESG-relatedness (+26% improvement) and 84% ESG action-orientation (+34% improvement), while capturing a high number of unique ESG terms 359 under the exact implementation; 2555 under the flexible implementation). Additionally, human evaluation confirms that ESGSenticNet correctly categorises concepts under their respective ESG topics with 86%+ accuracy. As a result, during topic analysis, the topic terms matched from ESGSenticNet align with their associated ESG topics with considerable reliability. Finally, as a lexical method, ESGSenticNet requires no training or technical expertise, making it accessible to stakeholders across diverse backgrounds. To underscore the significance of these results for ESG topic analysis, we relate them back to the three core challenges—irrelevance, immateriality, and limited organisation—as well as the practical concerns of comprehensiveness & diversity and computational complexity. While the three core challenges have been previously defined, comprehensiveness & diversity captures the extent and variety of ESG content each method is able to extract, and computational complexity refers to the ease of implementation and resource demands of each method. Table 1 qualitatively compares ESGSenticNet, topic models, and existing lexicons in terms of how well they address each of these five dimensions. These assessments are informed by our experimental results—ESG-relatedness corresponds to irrelevance, ESG action-orientation to immateriality, and the number of unique ESG terms to comprehensiveness & diversity—as well as qualitative evaluations for limited organisation and computational complexity.
[…]
Topic Models Extracted terms are largely irrelevant to ESG topics Extracted terms rarely reflect concrete ESG action-orientation, reflecting immateriality Unstructured and not aligned with ESG frameworks Specific models capture a high amount of ESG terms, although majority of extracted terms are not ESG related High: requires extensive hyperparameter tuning and model selection
Existing Lexicons More relevant than topic models, but still often include unrelated terms Partially captures but still largely omits ESG action-oriented terms Structured by a single ESG taxonomy Captures a small to moderate amount of ESG terms Low: no training or tuning required
[…]
Table 1: Qualitative comparison of ESG text analysis methods across five key dimensions. Lim-Org stands for limited organisation, Compre-Div for comprehensiveness & diversity, Complexity for computational complexity. Each assessment reflects evidence from our experimental findings where applicable — for example, high irrelevance corresponds to low ESG-relatedness scores, high immateriality to low ESG action-orientation scores, and limited comprehensiveness to low unique ESG terms captured. Limited organisation and computational complexity are assessed through qualitative evaluations.
The findings position ESGSenticNet as a tool that significantly enhances ESG topic analysis, offering clear advantages over existing approaches across all five dimensions.
The rest of our paper is arranged as follows. We begin with a literature review (Section 5). Then, we explain the preliminary steps (Section 4) prior to constructing the knowledge base, which involves describing the data utilised in this study (Section 4.1), and defining the knowledge base’s organisational framework. The latter involves establishing a hierarchical ESG taxonomy that defines the topics in our knowledge base (sections 4.24.3) and specifying the relations and rules that govern its structure (Section 4.4).
After outlining the prelimnary steps, we delve into our neurosymbolic framework for constructing the knowledge triples within ESGSenticNet (Section 4.5). This involves the extraction of concepts via ESG concept parser (Section 4.6), quality control—processing the extracted concepts to ensure their coherence (Section 4.7), followed by the labelling process for classifying these concepts according to our taxonomy topics (Section 4.8). This labelling process involves leveraging GPT-4o inference, supported by a graph-based semi-supervised method (sections 4.8.14.8.24.8.34.8.4).
Our neurosymbolic framework (Section 4.5) results in the construction of ESGSenticNet, which we extensively evaluate through experiments on topic analysis, with the results and implications discussed (Section 2).
2 Results
To evaluate the effectiveness of ESGSenticNet against existing NLP methods, we perform topic analysis on the evaluation set (described in Section 4.1). This set comprises 319 sustainability reports from SGX companies, and the reports were unused during the development of ESGSenticNet. In the following, we discuss the evaluation metrics (Section 2.1) and experimental settings (Section 2.2) utilised during this experiment, before discussing the experimental results (Section 2.3). Thereafter, we further evaluate the accuracy of ESGSenticNet’s triples through human evaluation (Section 2.4), before discussing a case study of ESGSenticNet’s application (Section 3.1) and the implications of our ESGSenticNet study (Section 3.2). We finish this section with additional ablation studies (Section 4.10) to improve the robustness of our findings.
2.1 Evaluation Metrics
As discussed earlier (Section 1), ESG topic analysis has distinct requirements as it supports sustainability assessments. It focuses on extracting ESG content that is significant for these assessments, and organised within established ESG frameworks. Therefore, while topic analysis metrics—specifically pertaining to topic models (i.e. perplexity, PMI) [1]—have been explored, ESG topic analysis require specialised metrics. Therefore, methods are evaluated according to the following:
  • ESG relatedness: The proportion of all topic terms that are ESG-related, highlighting the extent to which yielded topics are relevant to ESG and therefore ESG analysis.
  • ESG action orientation: The proportion of all topic terms that express an action taken toward ESG, or in other words, an ESG effort. This indicates whether the terms are able to express information material to ESG assessments, and of interest to stakeholders.
For the above metrics, GPT-4 is leveraged to classify all topic terms yielded from the models, aligning with the LLM-as-a-judge method from established NLP studies [80]. GPT-4 distinguishes the following – 1) whether each topic term is ESG-related, 2) whether each topic term constitutes an action toward improving ESG. Thereafter, we aggregate the GPT-4 classification results to produce scores for ESG relation, unique ESG terms, ESG action. We utilise GPT-4 in our work given that it offers a scalable, automated, and reproducible approach to evaluating ESG topic terms [18] — a task that would be prohibitively costly and impractical to perform manually, given the tens of thousands of topic terms involved. At the same time, we recognise the potential limitations of LLM-based evaluation [80]. Therefore, to ensure robustness, we validate the use of GPT-4 in our evaluation through an GPT4-human agreement study (Section 4.10.2). Full experimental details for topic analysis evaluation are described in Section 2.2, with the prompts provided in Figure 1 (ESG relatedness), and Figure 2 (ESG action orientation).
[Task Description] As a sustainability expert, your task is to classify terms based on their relevance to Environmental, Social, and Governance (ESG) factors. [Response Instructions] Evaluate each term based on its content to determine its most appropriate topic. If a term is explicitly related to ESG, assign the label ‘ESG’. If a term is not related to ESG or the relationship is vague, assign the label ‘non-ESG’. Ensure all responses strictly adhere to the definitions and format provided. [Response Format] Fill in your response in the triple backticks below using the following format: Input: <term> Response: <ESG or non-ESG>
Figure 1: Prompt for evaluating ESG relatedness
Figure 2: Prompt for evaluating ESG action orientation
2.2 Experimental Settings
For this study, the following baselines were utilised. Topic models that are algebraic – NMF [40], probabilistic – LDA [10], neural – ProdLDA [67], TSCTM [74], BERTopic [30], ECRTM [73], and other publicly available ESG dictionaries – Baier [5] Naiara [60]. To the best of our knowledge, ESG dictionary methods in the published literature are limited, with Baier and Naiara among the only few works that make their full dictionaries publicly available.
[…]
Dictionary methods comprise terms that are categorised according to ESG topics, with a topic’s occurrence computed from the frequency of occurrence of its associated terms. In our experiment, the topic terms of the dictionaries correspond to dictionary terms that appear within the corpora of sustainability reports.
ESGSenticNet comprises knowledge triples in the format (concept, relation, ESG topic), delineating how specific concepts related with ESG topics or topics. To enhance semantic specificity and context, we employ three word concepts, exclusively using those that possess the ‘supports’ relation toward a ESG topic. Similar to dictionary methods, a topic’s occurrence is determined by the frequency of appearance of concepts that support the topic (we explore this further in a case study in section (3.1)). The topic terms of ESGSenticNet correspond to these concepts that appear within the sustainability reports corpora. Additionally, we experiment with two configurations for ESGSenticNet: Exact involves exact string matches of each concept, Flexible involves taking advantage of the (verb, noun phrase) arrangement of concepts for matching. To elaborate on this, the verb (1st word) is matched separately from the corresponding noun-phrase (2nd & 3rd words) within each sentence. Given the example “We reduce our water consumption”, ‘reduce’ is matched as the verb, and ‘water consumption’ is matched as the noun-phrase, with ‘reduce water consumption’ returned as the matched concept without requiring its exact string match. Flexible enables greater detection of ESG concepts by allowing for semantic variability (see study in Section 4.10.3), while Exact is useful for detecting precise ESG impacts.
2.3 Experimental Results
Baseline Best ESG-unique Best ESG-rel Best ESG-act
NMF 332* 0.28* 0.08* 3 0.42 0.00 332* 0.28* 0.08*
LDA 23* 0.31* 0.01* 3 0.36 0.00 23* 0.31* 0.01*
ProdLDA 334* 0.26* 0.07* 17 0.38 0.10 13 0.38 0.16
BERTopic 76 0.33 0.05 13 0.40 0.04 13 0.35 0.05
TSCTM 359 0.15 0.06 11 0.32 0.06 23 0.31 0.15
ECRTM 440 0.18 0.09 23 0.46 0.08 182 0.31 0.15
Baier 212* 0.49* 0.29* 212* 0.49* 0.29* 212* 0.49* 0.29*
Naiara 1* 0.50* 0.50* 1* 0.50* 0.50* 1* 0.50* 0.50*
Figure 3: Heatmap highlighting performance of different topic analysis methods on the metrics of number of unique ESG terms (ESG-unique), ESG relatedness (ESG-rel), ESG action orientation (ESG-act). ESGSenticNet(Exact) refers to ESGSenticNet deployed by Exact matching, while ESGSenticNet(Flexible) refers to ESGSenticNet deployed by Flexible matching, as explained in Section 5.2. Darker green indicates better performance, lighter green indicates poorer performance.
Our experimental results (Table 2) indicate that ESGSenticNet scores higher than the other topic analysis methods on ESG relatedness and ESG action orientation by at least 26% and 31% respectively. Additionally, both ESGSenticNet (exact) and ESGSenticNet (flexible) extract a high number of unique ESG terms, with ESGSenticNet (flexible) scoring the highest amongst all methods tested. Although requiring limited computational resources as a lexical method, ESGSenticNet produces significantly higher performance compared to the more expensive unsupervised topic modeling methods. This is despite how the topic models were exhaustively ran on numerous permutations of parameters (Section 2.2) that required extensive computational costs.
Method Topic Example Topic Words
Topic 2 rsp, venture, johor, black, sites
Topic 2 banks, grand employees, duty, palm
ProdLDA Topic 1 financial, value, non assets, december
Topic 2 loss, company, board, statements, directors
NMF Topic 1 duly, assets, transparency, value, priority
Topic 2 property, fair, rate, million, new
Table 3: Examples of topic words for topic models with the highest scores on the number of unique ESG terms.
Unsupervised topic models such as TSCTM, ECRTM, ProdLDA, NMF can extract a high number of unique ESG terms (>300>300). Yet, for the same runs, the concentration or prevalence of ESG terms within each topic is modest, as highlighted by their limited scores in ESG relatedness (<0.30<0.30). Therefore, despite capturing diverse ESG topic terms, these methods fall short in generating topics that are richly centered on ESG themes. From Table 3, we highlight examples of the topics generated by these methods in comparison to ESGSenticNet, highlighting how most of the topics generated by these topic models are not ESG-centric. This limitation is critical in ESG topic analysis, given that the primary goal is to derive insights from topics that are not only inclusive of ESG terms but are also substantially focused on ESG issues.
We also compare the number of unique ESG-related terms extracted by methods that construct structured ESG lexicons from existing sustainability corpora—specifically Baier and Naiara, as this serves as an indicator of lexical coverage and diversity. ESGSenticNet yields 359 unique terms in its exact matching configuration and 2,555 under its more flexible variant, compared to 212 for Baier and just 1 for Naiara. These differences suggest that ESGSenticNet captures a broader and more diverse set of ESG terms from the test corpora, reflecting a more extensive and diverse collection of ESG lexicons (or ESG concepts) relative to the Baier and Naiara lexicons. This outcome reflects the advantages of our neuro-symbolic framework—discussed in Section 5.2—over manual or anchor-based lexicon construction methods, as it enables scalable and systematic identification of ESG concepts through automated concept parsing, GPT-4o-based annotation, diversity-aware seed selection, and label propagation.
Moreover, topic model runs that yield the highest ESG relatedness are only able to yield few unique ESG-terms (<25<25). Furthermore, for these topic model configurations, as well as for other dictionaries such as Baier and Naiara, ESG relatedness is still $\leq 0.50$. This suggests that yielded topics are still not predominantly centred around ESG themes.
Finally, most of the tested baseline methods score modestly for ESG action orientation, Naiara noticeably higher. Yet, Naiara is only able to derive $1$ unique ESG-term, which limits the significance of its analysis. These results suggest that the baseline methods may not fully capture information that is material to stakeholders, particularly in terms of conveying ESG actions and efforts.
2.4 ESGSenticNet Human Evaluation
In this section, we evaluate ESGSenticNet’s effectiveness in categorising concepts under the correct topics. This ensures that during topic analysis, the topic terms matched from ESGSenticNet reliably correspond to their intended ESG topics. To assess this, human annotators determine the accuracy of relations between concepts and topics. 500 concepts are randomly selected from our knowledge base. A concept is found within a knowledge triple, possessing a relation with a topic in the format of (concept, relation, ESG topic). A correct relation represents that the concept is accurately classified under the ESG topic. The accuracy of the relations are evaluated separately for each topic type—pillar, broad, cross-broad, sub, cross-sub, with each of them containing different ESG topics as described in Table 7. Human annotators adhere to the following instructions (B provides more details on our annotators and annotation scheme):
The table below presents various concepts along with their polarities and respective relations toward ESG topics. Each concept may be associated with multiple relations across different ESG topics. These topics are classified into five topic types: pillar, broad, sub, cross-broad, and cross-sub. If a concept has no relevant association with a topic, it is marked as ‘not applicable’. Your task is to evaluate whether each concept is correctly related to its corresponding topics. For each topic, determine if the relation is the most appropriate compared to the other topics within the same topic type. Mark ‘TRUE’ if the relation and topic are correct, and ‘FALSE’ if they are incorrect. Additionally, assess the accuracy of the polarity of each concept, and mark ‘TRUE’ if it is correct and ‘FALSE’ if it is incorrect.
Topic Type Accuracy
Pillar 95.0%
Broad 89.3%
Cross-Broad 86.1%
Sub 87.8%
Cross-Sub 90.1%
Table 4: Accuracy of ESGSenticNet for the relations across all topic types.
Results from the human evaluation study highlight that ESGSenticNet retains a high degree of accuracy at 86+% for relations across all topic types, as shown in Table 4. These findings suggest that ESGSenticNet has reasonable accuracy in categorising concepts under the correct topics. Consequently, the topic terms matched from ESGSenticNet during topic analysis corresponds to its related ESG topic with considerable reliability.
3 Discussion
3.1 Case Study - Insights offered by ESGSenticNet
Figure 4: Frequency of ESGSenticNet topics from the flexible run
The topic analysis conducted by ESGSenticNet in Section 2 is visualised with the Figure 4 & Figure 5, which represent different ways to analyse the text corpora. Figure 4 provides a higher level breakdown of the topics present in the corpora. It highlights the top 5 frequently occurring broad ESG topics present in the corpora, as well as the proportion of their constituent sub topics, in accordance with ESGSenticNet’s hierarchical taxonomy. The topic frequencies are denoted by the occurrence of matched concepts that ‘support’ a topic, leveraging ESGSenticNet’s predefined relations. Separately, Figure 5 provides a more granular analysis of the corpora, highlighting matched concepts that explicitly convey ESG actions, thereby communicating important sustainability information. Figure 5 displays the top occurring concepts that support the topics shown in Figure 4. Put together, Figure 4 and Figure 5 highlight different methods to analyse a sustainability corpus, which can be utilised depending on a stakeholder’s interest and preference.
While the case study presents insights derived from a test corpus of 319 SGX sustainability reports, the same form of topical analysis can be applied across different corpora of sustainability reports in various real-world scenarios. We detail examples of these in the following (to qualify, real-world use cases may extend beyond those listed):
  • Analysts for ESG funds can leverage ESGSenticNet’s topical insights to evaluate the sustainability of corporations, which is useful for sustainable portfolio construction [35]. In this context, where timely, transparent, and interpretable assessments are essential, using an on-demand and open-source tool like ESGSenticNet for rule-based topic analysis is advantageous over relying on ESG scores, which are often obscure, inconsistent, and infrequently updated [6, 32].
  • Regulators can utilise ESGSenticNet to perform sector-wide audits and benchmark companies’ sustainability disclosures against each other. For instance, by applying ESGSenticNet across disclosures from firms within a particular industry, regulators can quickly identify under-reporting on material ESG issues (e.g., labour safety in the construction sector, emissions control in the logistics sector) [27]. This capability enables data-driven supervision and targeted follow-ups. Moreover, as ESGSenticNet operates as an interpretable, rule-based method, it provides transparency in how disclosures are assessed—crucial for public accountability and trust in regulatory oversight.
Additionally, compared to the other ESG topic analysis methods (i.e. existing dictionaries and topic modeling), ESGSenticNet’s deployment holds significant advantages in that it extracts – 1) More relevant ESG insights, as demonstrated by the greater proportion of ESG related terms (highest ESG-relatedness score in Table 7). 2) ESG insights with greater materiality for stakeholders (highest ESG-action orientation score in Table 7). 3) ESG insights that are organised according to established ESG frameworks (i.e. the extracted topic terms from ESGSenticNet encompass ESG topics that are structured according to established ESG themes (e.g., compliance, management), as shown in Figure 12, due to ESGSenticNet’s taxonomy centric organisation (Section 3).). These advantages position ESGSenticNet as a topic analysis tool that could potentially be more effective than existing solutions for the aforementioned scenarios.
3.2 Implications
Our results indicate that ESGSenticNet outperforms existing methods in capturing a significantly higher proportion of ESG-related and ESG-action oriented topic terms while also identifying a large number of unique ESG terms. This suggests that stakeholders using ESGSenticNet for ESG topic analysis can extract more comprehensive and relevant insights from sustainability reports than existing approaches, thereby enhancing ESG assessments. Crucially, this capability could enable new discoveries in sustainable finance and sustainability compliance research, such as examining how textual patterns in sustainability reports relate to firm value [71] or sustainability violations [57]. ESGSenticNet facilitates structured information retrieval by identifying ESG topics—organised according to a hierarchical taxonomy—within a report and surfacing the specific concepts that constitute them. This structured output can also support document classification by enabling stakeholders to flag disclosures with distinctive reporting patterns, such as under-reporting in key sustainability areas [27], as discussed in Section 3.1. These capabilities collectively support real-world decision-making use cases, including portfolio construction for ESG fund managers [35] and sector-level audits for regulators [27], as outlined in Section 3.1.
Moreover, our findings highlight that ESGSenticNet can function effectively as a structured lexicon without requiring additional training or tuning, while producing superior results over expensive state-of-the-art NLP topic models. This can make ESG topic analysis more accessible to stakeholders with limited technical expertise or computational resources. Given the interdisciplinary nature of sustainability research, this accessibility is essential for enabling broader adoption by researchers across different fields, thereby facilitating diverse perspectives and approaches that can drive further discoveries in sustainability research.
Additionally, ESGSenticNet’s superior performance over existing methods highlights the effectiveness of our scalable and automated neurosymbolic AI approach for lexicon construction in ESG topic analysis. Given its effectiveness, this framework could also be adapted to expand lexicons in other key sustainability contexts that require topic analysis—such as climate risk disclosures [49] and environmental policy analysis [12]—where large-scale lexicon development is constrained by high human annotation costs, making automation a necessity.
Moreover, this scalable and automated framework also allows the knowledge base to be continuously updated at minimal cost, ensuring that the lexicons remain relevant over time. This is particularly important in the ESG domain, where sustainability reporting content is dynamic and constantly evolving [22]. As sustainability reports change, maintaining up-to-date lexicons is essential for accurately capturing emerging ESG themes and patterns.
4 Methods
In this section, we introduce the data utilised in this study (Section 4.1), before defining how ESGSenticNet’s triples are organised, laying the groundwork for their subsequent derivation in Section 4.5. As explained earlier in the introduction (Section 1), ESGSenticNet comprises knowledge triples in the format of (concept, relation, ESG topic). The ESG topics within these triples are drawn from our hierarchical taxonomy in Section 4.3, whose construction methodology is detailed in Section 4.2. Additionally, we define the relations between concepts and ESG topics within the triples, along with the rules governing these relations, in Section 4.4.
4.1 Data
This study utilises a total of 1,998 sustainability reports from 589 Singapore Exchange (SGX) based companies, in the period of 2015 to 2023 inclusive. We randomly divide all the reports into the construction set (1,679 reports) and the evaluation set (319 reports) with a ratio of 85:15, without any overlap between the two sets. The construction set is utilised for developing ESGSenticNet, enabling taxonomy derivation (Section 4.2) and knowledge base construction (Section 4.5). The evaluation set is reserved for evaluating ESGSenticNet (Section 2). All reports are publicly accessible online, and are downloaded in PDF format (via the respective company and company-affiliated websites). Thereafter, we processed the report into text format via the PyPDF2 tool. For reproducibility, we provide the full details on the reports in the supplementary material and instructions on how to access these reports. Specifically, we provide CSV files that specify the companies and the corresponding years of each report. Using these details (company name & report year), all sustainability reports can be easily searched for and downloaded online. Due to space constraints, we can only provide a sample list of the reports we obtained below in Table 5.
Company Report Date
ASIAPHOS LIMITED 2019-12
MEGROUP LIMITED 2020-03
LHN LIMITED 2020-09
CHINA REAL ESTATE GRP LIMITED 2018-06
CITIC ENVIROTECH LIMITED 2018-12
JAWALA INC. 2019-07
BUND CENTER INVESTMENT LIMITED 2018-12
SAMURAI 2K AEROSOL LIMITED 2021-03
HOR KEW CORPORATION LIMITED 2019-12
HALCYON AGRI CORPORATION LIMITED 2021-12
PROPNEX LIMITED 2019-12
NAM CHEONG LIMITED 2018-12
HC SURGICAL SPECIALISTS LIMITED 2019-05
GKE CORPORATION LIMITED 2019-05
Table 5: Sample list of SGX sustainability reports provided in the supplementary material.
4.2 Taxonomy Construction Methodology
[…]